# From Unitarity to Irreversibility: The Role of Infinite Tensor Products and Nested Wigner’s Friends

### v1

@@ I Introduction
In the von Neumann scheme for ideal quantum measurement [1, 2], the ‘object’ is prepared in an initial state $|\psi\rangle$. With respect to a ‘mismatching’ context (relative to that preparation)—or, equivalently, orthonormal basis or maximal operator [3, Satz 8] ([4, Theorem 1, § 84])—$|\psi\rangle$ is in a coherent superposition (linear combination) $|\psi\rangle=\sum_{i=1}^{n}a_{i}|\psi_{i}\rangle$ of (basis) elements $|\psi_{i}\rangle$ of that context.
The ‘measurement ancilla’—along with synonymous terms such as ‘provision’, ‘component’, or ‘arrangement’—should be represented by another state, denoted as $|\varphi\rangle$. This state, in relation to a suitable basis ${|\varphi_{1}\rangle,\ldots,|\varphi_{n}\rangle}$, can also be expressed as a coherent superposition: $|\varphi\rangle=\sum_{j=1}^{n}b_{j}|\varphi_{j}\rangle$. When an interaction occurs between the ‘object’ and the ‘measurement ancilla’, the combined state $|\Psi\rangle$ becomes a non-factorizable tensor product, meaning $c_{ij}\neq a_{i}b_{j}$.
$|\Psi\rangle=\sum_{i,j=1}^{n}c_{ij}|\psi_{i}\rangle\otimes|\varphi_{j}\rangle=\sum_{i,j=1}^{n}c_{ij}|\psi_{i}\varphi_{j}\rangle.$ (1)
@@ II.1 Elementary Tensors as Products of Basis Vectors
[…]
@@ II.2 Tensor Product Space
    We define the inner product on elementary tensors by the product of the individual inner products [7, Definition II.5., p. 63]:
@@
By defining elementary tensors as products of basis vectors from the bases of the factors, a concrete and manageable set of elementary tensors is obtained, spanning the tensor product space. This approach, derived from finite tensor products [4, Theorem 1, § 24,25], simplifies both the definition and the computation of the inner product, ensuring that the space $\bigotimes_{n=1}^{\infty}\mathcal{H}_{n}$ has a well-defined Hilbert space structure.
However, this construction does not directly address the uncountable infinity of elementary products. To illustrate this, we can draw an analogy with the representation of real numbers as expansions in an $n$-ary system, where they are encoded using a (finite) set of basis elements. Just as Cantor’s diagonal argument shows that the reals cannot be enumerated by any countable set of indices, so too can we not enumerate the uncountable infinity of elementary products in the infinite tensor product. In von Neumann’s own words “generalisations of the direct product lead to higher set-theoretical powers (G. Cantor’s ”Alephs” )” [8, S 4, p. 4]. Following Von Neumann’s ‘incomplete infinite direct products’ [8, Chapter 4], Thirring and Wehrl define the infinite tensor product in terms of equivalence classes [9, § 2] (see also Thirring [7, Definition II.4., p. 63]) discussed later in the context of sectorization. One could even go so far as to suspect that many of the upcoming issues related to continuity originate from this fact.
@@ II.3.3 Bounded Operators
For instance, consider the vector $|\psi\rangle=|\uparrow\rangle\otimes|\uparrow\rangle\otimes\cdots$, where $F|\psi\rangle=|\psi\rangle$ and $\|F|\psi\rangle\|=\||\psi\rangle\|=1$. On the other hand, for any vector containing a component orthogonal to $\begin{pmatrix}1,0\end{pmatrix}^{\intercal}$ in at least one factor $|\downarrow\rangle=\begin{pmatrix}0,1\end{pmatrix}^{\intercal}$, $F$ maps it to the zero vector. For example, for the vector $|\varphi\rangle=|\uparrow\rangle\otimes|\uparrow\rangle\otimes\cdots\otimes|\downarrow\rangle\otimes|\uparrow\rangle\otimes\cdots,$ we have $F|\varphi\rangle=0$.
@@ II.4 Sectorization
Von Neumann’s concept of ‘incomplete infinite direct products’ [8, Chapter 4], as reflected in the notion of superselection sectors in algebraic quantum field theory [11, 12, 13, 14, 15], provides a solution to the problem that a single (or finitely many) subfactor(s) could nullify the inner product by ‘grouping’ vectors that differ from each other in only finitely many subfactors, or are otherwise ‘close to’ each other. These groupings are mutually orthogonal and can be demonstrated to be equivalence classes referred to as sectors. This implies that vectors from different sectors differ in an infinite number of subfactors and are orthogonal in the sense that their scalar product is zero. In von Neumann’s own words, “What happens could be described in the quantum-mechanical terminology as a ‘splitting up’ of [[the full tensor product]] into ‘non-intercombining systems of states’, corresponding to the ‘incomplete’ direct products” [8, § 6, p. 4]
Formally, within each sector are only those infinite tensor products that are located ‘close to’ each other, such that their deviations from each other are ‘small’. Two vectors $|\Psi\rangle$ and $|\Phi\rangle$ are in the same sector if they are equivalent, denoted by $|\Psi\rangle\sim|\Phi\rangle$, when all but finitely many subfactors are either equal or unitary equivalent and to or ‘close to’ one another (only a unitary transformation apart); that is, with the notation from Eq. (2) we require [10, Eq. (9)]
@@
Therefore, for infinite tensor products, instead of directly dealing with the entire infinite tensor product space, one should consider regions or sectors within it. These sectors are equivalence classes of vectors that differ only by a finite number of components in the tensor product, or are otherwise close to (unitary equivalent) each other. These sectors represent different ‘global’ or ‘macroscopic’ configurations of the system [16].
[…]
@@ II.5 Factorization
It is reasonable to identify (orthogonal) projections of type In factors with elements of orthonormal bases, or equivalently, with contexts, blocks in quantum logic, or maximal operators. This identification is supported by [3, Satz 8] (see also [4, § 82]). Type In factors are the only ones in finite dimensional Hilbert space. They have minimal orthogonal projections (self-adjoint and idempotent) that correspond to one-dimensional subspaces of the Hilbert space, as well as convex combinations thereof (projecting into higher-dimensional subspaces). Indeed, any sequence of mutually orthogonal (orthogonal) projections $|\psi_{i}\rangle$, combined with any sequence of probabilities $p_{i}\in[0,1]$ satisfying $\sum_{i=1}^{k}p_{i}=1$ where $k\leq n$, forms a density operator $\rho=\sum_{i=1}^{k}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$.
@@
Despite being diffuse, the trace of a projection in a type II factor is still faithful, normal, and semi-finite. A faithful trace is one that does not vanish on any non-zero positive element of the von Neumann algebra. A normal trace respects the convergence of operators (in the weak topology), ensuring that the trace of a limit of operators equals the limit of their traces. A semi-finite trace is one such that for any positive element, there is a non-zero ‘sub-element’ on which the trace is finite. For type II1 factors, the trace assigns a value in the continuous interval $[0,1]$, where 0 corresponds to the zero projection and 1 corresponds to the identity projection. This trace function behaves like a measure of ‘dimension’ but is not tied to integer dimensions as in finite-dimensional spaces. For type II factors, the trace can take values in $[0,\infty]$.
@@
The origin of the term ‘factor’ may come from a tensor product factorization: Suppose $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Then $F_{1}=\mathcal{B}(\mathcal{H}_{1})\otimes\mathbb{1}$ and $F_{2}=\mathbb{1}\otimes\mathcal{B}(\mathcal{H}_{2})$ are factors of $\mathcal{B}(\mathcal{H})$ [21, Exercise 3.3.8].
@@ III Nested Wigner’s friends as infinite tensor products
Grangier and Van Den Bossche have recently proposed that the apparent loss of coherence in such situations is attributable to sectorization and the consequent loss of unitary equivalence within finite systems [26, 10, 27, 28], as previously discussed in subsection II.4. According to their proposal, sectorization is a physical process in infinite algebras: Separable sectors correspond to ‘classical outcomes’ and ‘macroscopic states’ of pointers [16, 29, 30]. Within a sector, the context and the resulting measurement have fixed macroscopic values, but a finite (sub)system within each sector can still undergo unitary evolution as long as the state remains within its sector. Any new measurement results in a shift to a different sector, thus establishing a new context and corresponding outcome. In the extreme case, a sector can be associated with a single element of a context. Contexts and their associated sectors are ‘reshuffled’ or ‘scrambled’ when new incompatible measurements are performed.
@@
For type I factors, some of these operators can be interpreted as pure (that is, ‘minimal’ one-dimensional orthogonal projection operators) or mixed states. Indeed, intuition from finite tensor products suggests that, through spectral decomposition of the operators in a factor, the respective orthogonal projections in type In factors correspond to elements of contexts (associated with maximal operators, see [3, Satz 8]). This implies that they refer to pure states spanned by vectors in the respective Hilbert space.
@@
More explicitly, as has been pointed out earlier in the context of difficulties in defining the inner product, any slight mismatch between (successive) friends’ measurements ‘builds up’ into a total loss of coherence. This results in a vanishing inner product $\langle\Psi|\Psi^{\prime}\rangle$ which converges to zero, indicating that the product states $|\Psi\rangle$ and $|\Psi^{\prime}\rangle$ are orthogonal even if each single mismatch characterized by $\langle\psi_{i}|\psi_{i}^{\prime}\rangle$ is very close to 1. This type of ‘decoherence’ is gradual and smooth in the sense that there is no abrupt discontinuous transition—indicating a well-defined, localizable Heisenberg cut at some scale—but a gradual, continuous loss of information about the initial state: Let $10\ll|\langle\psi_{i}|\psi_{i}^{\prime}\rangle|=\delta_{i}<$ 1 be this match per friends $i$ and $i^{\prime}$, then
@@
Another approach to addressing Bell’s argument involves the concept of infinity processes, as discussed by Weyl [34, pp.41,42] in the context of Zeno’s paradoxes (of infinite divisibility). If we consider the continuum (or at least the infinite divisibility of space and time), “if the segment of length 1 really consists of infinitely many subsegments of lengths $1/2,1/4,1/8,\ldots$, as of ‘chopped-off’ wholes, then it is incompatible with the character of the infinite as the ‘incompletable’ that Achilles should have been able to traverse them all.” This implies that even for classical motion in a continuum to be possible, we require transfinite capacities. This concept can be applied to the infinite nesting of Wigner’s friends by considering this nesting or chaining as the effective oneness which we experience; resulting in irreversible measurements in the transfinite limit.
@@ IV Historical Analogues
This section explores several related but distinct concepts that have been investigated in various areas of physics.
@@ V Summary
Resolutions of the quantum measurement problem and the Umkehreinwand in statistical physics through means relativity involve epistemological commitments. Previous efforts to simulate measurement processes with von Neumann algebras, like those of Hepp [16], have been questioned for using transfinite concepts without operational validation [33, 30].
@@
Yet, we must remain cognizant that both FAPP and transfinite irreversibility remain mathematical constructs, ‘images of our imagination’ [54] which ultimately are justified by their practical usefulness and correspondence with phenomenology. We should therefore exercise caution, not conflating the practical utility of our models with absolute certainty about physical reality.

### v2

@@ I Introduction
Unitary equivalence can be surpassed through infinite means. This observation is consistent with findings in number theory and analysis, where finite operations on rational numbers cannot yield results beyond the rational domain. However, when infinite methods and techniques are employed—such as Cantor’s diagonalization [1, 2] or the construction of Specker sequences [3, 4, 5], exemplified by Chaitin’s halting probability [6, 7]—it becomes possible to conceptualize irrational, incomputable, and random (algorithmically incompressible) numbers.
The central problem this paper addresses is the quantum measurement problem: how can the non-unitary, irreversible measurement process (von Neumann’s “process 1”) emerge from the purely unitary and reversible evolution described by the Schrödinger equation (“process 2”)? Mainstream approaches to this problem include decoherence-based accounts, which explain the apparent collapse as a result of entanglement with the environment; objective collapse theories, which modify quantum dynamics to include a physical collapse mechanism; and Everettian or Many-Worlds interpretations, which posit that all outcomes occur in different branches of a universal wavefunction. This paper explores an alternative perspective, investigating the hypothesis that the transition to irreversibility is an emergent phenomenon that arises in the mathematical limit of infinitely complex systems. This limit is modeled using infinite tensor products, with the nested Wigner’s Friend scenario serving as a conceptual framework for such an infinite regression. The advantage of this formalism is its ability to mathematically break unitary equivalence without altering the fundamental quantum postulates for finite systems.
The unitary group, which formalizes quantum state evolution (excluding irreversible measurements and processes such as tracings), is, like all groups, inherently ‘hermetic’ by definition. In particular, its closedness under unitary transformations reflects a fundamental property of group theory, connected to mere permutations or one-to-one transformations of the identity element. Consequently, it seems that irreversibility cannot emerge from purely unitary evolution. To explore this further, let us revisit the historical arguments that gave rise to the conundrum posed by von Neumann and others.
[…]
The ‘measurement ancilla’—along with synonymous terms such as ‘provision’, ‘component’, or ‘arrangement’—should be represented by another state, denoted as $|\varphi\rangle$. This state, in relation to a suitable basis $|\varphi_{1}\rangle,\ldots,|\varphi_{n}\rangle$}, can also be expressed as a coherent superposition: $|\varphi\rangle=\sum_{j=1}^{n}b_{j}|\varphi_{j}\rangle$. When an interaction occurs between the ‘object’ and the ‘measurement ancilla’, the combined state $|\Psi\rangle$
$|\Psi\rangle=\sum_{i,j=1}^{n}c_{ij}|\psi_{i}\rangle\otimes|\varphi_{j}\rangle=\sum_{i,j=1}^{n}c_{ij}|\psi_{i}\varphi_{j}\rangle$ (1)
becomes a non-factorizable tensor product, meaning that the coefficients $c_{ij}$ cannot be written as products $a_{i}b_{j}$.
@@ II.1 Elementary Tensors as Products of Basis Vectors
Given a countable collection of Hilbert spaces $\Big\{\mathcal{H}_{n}\Big|\;1\leq k_{n}\leq d_{n}\in\mathbb{N}\Big\}$, let $\Big\{|k_{n}\rangle\Big|\;1\leq k_{n}\leq d_{n}\in\mathbb{N}\Big\}$ be an orthonormal basis for each $\mathcal{H}_{n}$, where the dimension $d_{n}$ of $\mathcal{H}_{n}$ could be a finite positive integer or countably infinite.
@@ II.2 Tensor Product Space
    We define the inner product on elementary tensors by the product of the individual inner products [14, Definition II.5., p. 63]:
@@
By defining elementary tensors as products of basis vectors from the bases of the factors, a concrete and manageable set of elementary tensors is obtained, spanning the tensor product space. This approach, derived from finite tensor products [11, Theorem 1, § 24,25], simplifies both the definition and the computation of the inner product, ensuring that the space $\bigotimes_{n=1}^{\infty}\mathcal{H}_{n}$ has a well-defined Hilbert space structure.
However, this construction does not directly address the uncountable infinity of elementary products. To illustrate this, we can draw an analogy with the representation of real numbers as expansions in an $n$-ary system, where they are encoded using a (finite) set of basis elements. Just as Cantor’s diagonal argument shows that the reals cannot be enumerated by any countable set of indices, so too can we not enumerate the uncountable infinity of elementary products in the infinite tensor product. In von Neumann’s own words “generalisations of the direct product lead to higher set-theoretical powers (G. Cantor’s “Alephs”)” [15, S 4, p. 4]. Following Von Neumann’s ‘incomplete infinite direct products’ [15, Chapter 4], Thirring and Wehrl define the infinite tensor product in terms of equivalence classes [16, § 2] (see also Thirring [14, Definition II.4., p. 63]) discussed later in the context of sectorization. One could even go so far as to suspect that many of the upcoming issues related to continuity originate from this fact.
@@ II.3.3 Bounded Operators
[…]
@@ II.4 Sectorization
Von Neumann’s concept of ‘incomplete infinite direct products’ [15, Chapter 4], as reflected in the notion of superselection sectors in algebraic quantum field theory [18, 19, 20, 21, 22], provides a solution to the problem that a single (or finitely many) subfactor(s) could nullify the inner product by ‘grouping’ vectors that differ from each other in only finitely many subfactors, or are otherwise ‘close to’ each other. These groupings are mutually orthogonal and can be demonstrated to be equivalence classes referred to as sectors. This implies that vectors from different sectors differ in an infinite number of subfactors and are orthogonal in the sense that their scalar product is zero. In von Neumann’s own words, “What happens could be described in the quantum-mechanical terminology as a ‘splitting up’ of [[the full tensor product]] into ‘non-intercombining systems of states’, corresponding to the ‘incomplete’ direct products” [15, § 6, p. 4]
Formally, within each sector are only those infinite tensor products that are located ‘close to’ each other, such that their deviations from each other are ‘small’. Two vectors $|\Psi\rangle$ and $|\Phi\rangle$ are in the same sector if they are equivalent, denoted by $|\Psi\rangle\sim|\Phi\rangle$, when all but finitely many subfactors are either equal or unitary equivalent and to or ‘close to’ one another (only a unitary transformation apart); that is, with the notation from Eq. (2) we require [17, Eq. (9)]
@@
Therefore, for infinite tensor products, instead of directly dealing with the entire infinite tensor product space, one should consider regions or sectors within it. These sectors are equivalence classes of vectors that differ only by a finite number of components in the tensor product, or are otherwise close to (unitary equivalent) each other. In this framework, it is postulated that these mathematically distinct sectors correspond physically to different ‘global’ or ‘macroscopic’ configurations of the system, such as pointer states of a measurement apparatus [23].
States in different sectors cannot be coherently superposed by finite (unitary) means. One may say that, with respect to these finite unitary means, ‘coherence is lost’. Hepp even went so far as to state that “leads to macroscopically different ‘pointer positions’ and to a rigorous ’reduction of the wave packet’ ” [23]. Let us demonstrate this with an example. Consider an infinite array of spin-$\frac{1}{2}$ particles, where each particle has a Hilbert space $\mathcal{H}=\mathbb{C}^{2}$, spanned by the states $|\uparrow\rangle$ (spin up) and $|\downarrow\rangle$ (spin down). The entire system is then described by the infinite tensor product of these spaces $\mathcal{H}_{\text{total}}=\bigotimes_{i=1}^{\infty}\mathbb{C}^{2}_{i}$.
@@ II.5 Factorization
It is reasonable to identify (orthogonal) projections of type In factors with elements of orthonormal bases, or equivalently, with contexts, blocks in quantum logic, or maximal operators. This identification is supported by [10, Satz 8] (see also [11, § 82]). Type In factors are the only ones in finite dimensional Hilbert space. They have minimal orthogonal projections (self-adjoint and idempotent) that correspond to one-dimensional subspaces of the Hilbert space, as well as convex combinations thereof (projecting into higher-dimensional subspaces). Indeed, any sequence of mutually orthogonal (orthogonal) projections $|\psi_{i}\rangle$, combined with any sequence of probabilities $p_{i}\in[0,1]$ satisfying $\sum_{i=1}^{k}p_{i}=1$ where $k\leq n$, forms a density operator $\rho=\sum_{i=1}^{k}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$.
@@
Despite being diffuse, the trace of a projection in a type II factor is still faithful, normal, and semi-finite. A faithful trace is one that does not vanish on any non-zero positive element of the von Neumann algebra. A normal trace respects the convergence of operators (in the weak topology), ensuring that the trace of a limit of operators equals the limit of their traces. A semi-finite trace is one such that for any positive element, there is a non-zero ‘sub-element’ on which the trace is finite. For type II1 factors, the trace assigns a value in the continuous interval $[0,1]$, where 0 corresponds to the zero projection and 1 corresponds to the identity projection. This trace function behaves like a measure of ‘dimension’ but is not tied to integer dimensions as in finite-dimensional spaces. For type II factors, the trace can take values in $0,\infty]$.
@@
The origin of the term ‘factor’ may come from a tensor product factorization: Suppose $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Then $F_{1}=\mathcal{B}(\mathcal{H}_{1})\otimes\mathbb{1}$ and $F_{2}=\mathbb{1}\otimes\mathcal{B}(\mathcal{H}_{2})$ are factors of $\mathcal{B}(\mathcal{H})$ [28, Exercise 3.3.8].
@@ III Nested Wigner’s friends as infinite tensor products
Grangier and Van Den Bossche have recently proposed that the apparent loss of coherence in such situations is attributable to sectorization and the consequent loss of unitary equivalence within finite systems [33, 17, 34, 35], as previously discussed in subsection II.4. According to their proposal, sectorization is a physical process in infinite algebras where separable sectors correspond to ‘classical outcomes’ and ‘macroscopic states’ of pointers [23, 36, 37]. The “context” is physically realized by the choice of a measurement basis, which is determined by the experimental arrangement. An “observation” then corresponds to an interaction that entangles the system with the measurement apparatus, projecting the state into this chosen basis.
While a single such interaction is a standard unitary evolution of the combined system, the formalism of infinite tensor products suggests that an infinite sequence of such interactions forces the total state into one of the newly formed sectors. This transition to a different, orthogonal sector is what is meant by a “shift” that establishes a new macroscopic outcome. A subsequent measurement in an incompatible basis (a new context) would similarly drive the system into yet another distinct sector, a process described here as ‘reshuffling’ or ‘scrambling’ of contexts.
@@
For type I factors, some of these operators can be interpreted as pure (that is, ‘minimal’ one-dimensional orthogonal projection operators) or mixed states. Indeed, intuition from finite tensor products suggests that, through spectral decomposition of the operators in a factor, the respective orthogonal projections in type In factors correspond to elements of contexts (associated with maximal operators, see [10, Satz 8]). This implies that they refer to pure states spanned by vectors in the respective Hilbert space.
@@
More explicitly, as has been pointed out earlier in the context of difficulties in defining the inner product, any slight mismatch between (successive) friends’ measurements ‘builds up’ into a total loss of coherence. This results in a vanishing inner product $\langle\Psi|\Psi^{\prime}\rangle$ which converges to zero, indicating that the product states $|\Psi\rangle$ and $|\Psi^{\prime}\rangle$ are orthogonal even if each single mismatch characterized by $\langle\psi_{i}|\psi_{i}^{\prime}\rangle$ is very close to 1. This type of ‘decoherence’ is gradual and smooth in the sense that there is no abrupt discontinuous transition—indicating a well-defined, localizable Heisenberg cut at some scale—but a gradual, continuous loss of information about the initial state: Let $0\ll|\langle\psi_{i}|\psi_{i}^{\prime}\rangle|=\delta_{i}<1$ be this match per friends $i$ and $i^{\prime}$, then
@@
Another approach to addressing Bell’s argument involves the concept of infinity processes, as discussed by Weyl [41, pp.41,42] in the context of Zeno’s paradoxes (of infinite divisibility). If we consider the continuum (or at least the infinite divisibility of space and time), “if the segment of length 1 really consists of infinitely many subsegments of lengths $1/2,1/4,1/8,\ldots$, as of ‘chopped-off’ wholes, then it is incompatible with the character of the infinite as the ‘incompletable’ that Achilles should have been able to traverse them all.” This implies that even for classical motion in a continuum to be possible, we require transfinite capacities. This concept can be applied to the infinite nesting of Wigner’s friends by considering this nesting or chaining as the effective oneness which we experience; resulting in irreversible measurements in the transfinite limit.
@@ IV Historical Analogues
To motivate the use of infinite limits as a tool for explaining emergent irreversibility, this section presents analogies from other fields of mathematics and physics where infinite processes lead to qualitatively new phenomena not present in their finite counterparts. These examples are intended as heuristic support to demonstrate a recurring structural parallel, rather than as direct physical evidence for the quantum mechanical argument. This section explores several related but distinct concepts that have been investigated in various areas of physics.
@@ V Summary
The proposed mechanism can be contrasted with other solutions to the measurement problem. Unlike objective collapse models, it does not postulate a new physical collapse mechanism. In contrast to standard Everettian interpretations, the sector structure implies that not all branches of the universal state are equivalent, as transitions between them are forbidden by finite unitary means. And while the model shares features with decoherence, which also relies on interaction with a large environment, the use of infinite tensor products provides a formal basis for a strict, rather than merely practical (FAPP), breakdown of unitary equivalence between macroscopic outcomes.
The framework presented here also shares significant conceptual ground with standard decoherence theory, yet it is distinguished by a crucial formal difference. Both approaches attribute the loss of quantum coherence to the system’s entanglement with a larger, more complex entity—be it a physical environment or, in this paper’s model, an infinite chain of observers. The selection of a preferred basis (the ’pointer basis’ in decoherence) is analogous to the formation of sectors corresponding to macroscopic outcomes. The primary divergence, however, lies in the nature of the resulting irreversibility. Standard decoherence describes a practical (FAPP) process within a finite, albeit large, system-environment composite. The global evolution remains unitary, and coherence is merely ’leaked’ into the environmental degrees of freedom, becoming locally inaccessible but never truly destroyed. In contrast, the mechanism proposed here, leveraging the mathematical properties of the infinite tensor product limit, describes a formal and absolute breakdown of unitary equivalence. The emergence of sectors is not a matter of information being difficult to retrieve; it is a structural feature of the infinite-dimensional Hilbert space where states in different sectors are mathematically non-interconvertible by finite unitary means. Thus, while decoherence explains the appearance of classicality in a fundamentally unitary world, this paper’s formalism offers a mathematical route to the emergence of genuine, non-unitary irreversibility in the thermodynamic limit.
@@
Resolutions of the quantum measurement problem and the Umkehreinwand in statistical physics through means relativity entail significant epistemological commitments. Previous attempts to simulate measurement processes using von Neumann algebras, such as those by Hepp [23], have faced criticism for relying on transfinite concepts without operational validation [40, 37]. Nonetheless, these findings could be reconciled by adopting the perspective that “(FAPP) Infinity (FAPP) Does It.”
@@
Another straightforward pragmatic approach could be considered: Since this paper employs infinite processes to dispel unitary equivalence, one could avoid the infinite limit by transcribing the discussion into the framework of ‘for all practical purposes’ (FAPP). For operationalists who prefer to avoid the use of strict limits, the term ‘limit’ can be substituted with ‘FAPP unboundedness’ or ‘too-large-to-handle,’ and the symbol $\infty$ can be replaced with $\infty_{\text{\tiny FAPP}}$.
Yet, we must remain cognizant that both FAPP and transfinite irreversibility remain mathematical constructs, Hertz’s ‘images of our imagination’ [58] which ultimately are justified by their practical usefulness and correspondence with phenomenology. We should therefore exercise caution, not conflating the practical utility of our models with absolute certainty about physical reality.